OpenStudy (anonymous):

10x^4y^2 ---------- -5x^3y^2

OpenStudy (anonymous):

thats is one problem

OpenStudy (anonymous):

\[\frac{a^b}{a^c} = a^{b-c}\]

OpenStudy (anonymous):

ok the is the outline for the problem so wat is next this problem is much larger

OpenStudy (anonymous):

cancel out the exponents

OpenStudy (anonymous):

Break it up into the product of many fractions, then simplify using that rule.. \[\frac{10x^4y^2}{-5x^3y^2} = \frac{10}{-5}\cdot \frac{x^4}{x^3} \cdot \frac{y^2}{y^2}\]

OpenStudy (anonymous):

so it -2x

OpenStudy (anonymous):

Yes.

OpenStudy (anonymous):

Nicely done.

OpenStudy (anonymous):

thanks